### v1

@@ 6.4 Cyclomatic Complexity and General Statistics
Median
CC
Std. Dev.
CC
GPT-4 1000 104.54 1 33 3.01 2.71 1.60
Mistral-7B 10000 75.02 1 33.0 3.85 3.33 2.22
Falcon-180B 72000 72.06 1 46 4.34 3.33 3.10
Llama2-13B 5000 73.35 1 34.0 4.29 3.50 2.77
Gemini-Pro 40000 98.25 0 154 4.52 3.25 4.05
Codellama-13B 12000 83.30 1 94 4.63 3.20 4.64
Gemma-7B 47000 66.53 1 109 5.03 3.33 5.23
GPT-3.5 78000 96.52 0 240.0 6.10 4.43 5.68
We compute each model’s cyclomatic complexity (CC) number by randomly selecting $1000$ samples, as illustrated in Figure 10. The results unveil both expected and unexpected findings. While the exact parameter size of GPT-4 remains unknown, this model consistently generates the highest-quality code in terms of CC. Moreover, GPT-4 demonstrates remarkable performance on various benchmarks, such as HumanEval, indicating its proficiency in efficiently solving tasks without generating overly complex code. However, contrary to the expectations, the analysis of Table  10 reveals that GPT-4 does not necessarily produce shorter or simpler code. It tends to generate the longest C programs and exhibits the highest verification unknown score, implying that the verification process for ESBMC takes longer in the case of GPT-4 samples. Despite this, it manages to keep the CC number low. Among the examined LLMs, GPT-3.5 and Gemma7B demonstrate the poorest performance regarding cyclomatic complexity (CC). Interestingly, despite being a smaller model, Mistral7B closely follows GPT-4 in code quality. GPT-4, Mistral-7B, Falcon-180B, and Llama2-13B consistently exhibit low CC numbers across all samples, positioning them as the top performers regarding average CC. Conversely, other models display high maximum CC values, with corresponding high standard deviations, indicating less consistent performance. It is important to note, that due to the different total sample size, certain metrics like Max CC should be treated with caution. The more samples are generated the higher the chance for programs to appear with high CC numbers. Still, these metrics can serve as a good indicator, as for example Falcon-180B never managed to produce a program with a higher CC number then 46, despite the sample size for this model was the second largest. We aim to expand the dataset and equalise the sample size across the various models, to be able to make an even comparison. Despite this, such metrics are a useful indicator of performance.
(a) GPT4.
(b) MISTRAL-7B.
(c) FALCON-180B.
(d) LLAMA2-13B.
(e) GEMINI-pro.
(f) CODELLAMA-13B.
(g) GEMMA-7B.
(h) GPT-3.5.
Figure 10: Cyclomatic complexity of different LLMs on 1000 randomly selected samples.
@@ 6.5 Vulnerability Ranking
The Tables 11,12,13, and 14 provide an overview of the top 10 vulnerabilities produced by each model. Important to note that since the number of samples created by each model is different, the individual vulnerability counts do not serve as a basis for comparison.
Buffer Overflow on scanf ranks among the top three vulnerabilities identified across every LLM model. The functions scanf, fscanf, and sscanf do not limit the input size to the size of their respective buffers. This oversight leads to a risk of buffer overflow, which could enable an attacker to execute arbitrary code or cause a crash. As earlier noted, this issue can be associated with several CWEs, including the CWE-676, CWE-20, and CWE-787. While buffer overflow is a type of out-of-bounds write, CWE-787 encompasses a wider array of vulnerabilities, where CWE-120 is focused specifically on the classic scenario of buffer overflow resulting from not checking input sizes during buffer copy operations. Note that this analysis does not provide an accurate count of every identified CWE, but rather on the vulnerabilities verifiably identified by ESBMC.
The results indicate that LLMs consistently produce these errors in a zero-shot setting. While more complex issues such as arithmetic overflows and array bounds violations require a deeper understanding of the programming scenario, issues related to scanf would be comparatively simpler to avoid. Nonetheless, every model tested consistently exhibits buffer overflow errors on scanf, although the frequency varies.
Table 11: Top 10 Vulnerabilities in GPT-3.5 and FALCON-180B.
@@
GPT-3.5
1 $\mathcal{BO}$ Buffer overflow on scanf 84,363 38.05%
2 $\mathcal{DF}$ Dereference failure: NULL pointer 58,889 26.56%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 20,887 9.42%
4 $\mathcal{DF}$ Dereference failure: forgotten memory 4,788 2.16%
5 $\mathcal{DF}$ Array bounds violated: lower bound 7,836 3.53%
6 $\mathcal{DF}$ Array bounds violated: upper bound 7,773 3.51%
7 $\mathcal{AO}$ Arithmetic overflow on sub 6,422 2.90%
8 $\mathcal{DF}$ Dereference failure: array bounds violated 6,701 3.02%
9 $\mathcal{AO}$ Arithmetic overflow on add 5,169 2.33%
10 $\mathcal{AO}$ Arithmetic overflow on mul 4,266 1.92%
@@
1 $\mathcal{BO}$ Buffer overflow on scanf 52,811 34.99%
2 $\mathcal{DF}$ Dereference failure: NULL pointer 44,957 29.79%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 15,822 10.48%
4 $\mathcal{DF}$ Dereference failure: forgotten memory 5,942 3.94%
5 $\mathcal{DF}$ Dereference failure: array bounds violated 4,675 3.10%
6 $\mathcal{DF}$ Array bounds violated: upper bound 4,266 2.83%
7 $\mathcal{AO}$ Arithmetic overflow on sub 3,697 2.45%
8 $\mathcal{DF}$ Array bounds violated: lower bound 3,566 2.36%
9 $\mathcal{AO}$ Arithmetic overflow on add 2,880 1.91%
10 $\mathcal{BO}$ Buffer overflow on fscanf 2,555 1.69%
Table 12: Top 10 Vulnerabilities in LLAMA2-13B, and GEMMA-7B.
Rank Category Violation Type Count Percentage
@@
1 $\mathcal{DF}$ Dereference failure: NULL pointer 4,649 54.12%
2 $\mathcal{BO}$ Buffer overflow on scanf 814 9.48%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 811 9.44%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 435 5.06%
5 $\mathcal{DF}$ Dereference failure: forgotten memory 389 4.53%
6 $\mathcal{AO}$ Arithmetic overflow on add 236 2.75%
7 $\mathcal{AO}$ Arithmetic overflow on mul 179 2.08%
8 $\mathcal{DF}$ Array bounds violated: upper bound 176 2.05%
9 $\mathcal{AO}$ Arithmetic overflow on sub 152 1.77%
10 $\mathcal{BO}$ Division by zero 125 1.46%
gemma7b
1 $\mathcal{DF}$ Dereference failure: NULL pointer 89,888 65.76%
2 $\mathcal{BO}$ Buffer overflow on scanf 19,097 13.97%
3 $\mathcal{DF}$ Dereference failure: forgotten memory 5,491 4.02%
4 $\mathcal{DF}$ Dereference failure: invalid pointer 3,929 2.87%
5 $\mathcal{DF}$ Array bounds violated: upper bound 3,408 2.49%
6 $\mathcal{DF}$ Array bounds violated: lower bound 2,828 2.07%
7 $\mathcal{AO}$ Arithmetic overflow on sub 2,094 1.53%
8 $\mathcal{DF}$ Dereference failure: array bounds violated 2,060 1.51%
9 $\mathcal{AO}$ Arithmetic overflow on floating-point ieee_mul 1,458 1.07%
10 $\mathcal{AO}$ Arithmetic overflow on add 1,415 1.04%
Dereference failure-related issues were collectively the most numerous category for every LLM, and NULL pointer dereference consistently ranks first or second in every model. The generated code often includes pointers, but these models may not always correctly manage pointer operations or accurately simulate real-world applications’ complex memory management behaviors, leading to frequent dereference issues. However, LLMs, which operate primarily on pattern recognition without understanding underlying principles, often generate code with flawed pointer handling.
In addition, the datasets the were used for the training of these models often include a wide variety of pointer usage examples, which may not always adhere to safe programming standards. This, coupled with the complexities of dynamic memory management in languages like C, poses challenges for LLMs in consistently generating secure code. The recurrent issues with pointer dereferencing underscore the risks associated with deploying LLM-generated code in critical systems where security and reliability are paramount.
To mitigate these risks, it is crucial to develop enhanced training methodologies focused on robust memory handling. It is also paramount to implement advanced code analysis tools and frameworks. These tools will help detect and rectify vulnerabilities before deployment, thereby improving the security and reliability of LLM-generated code for real-world applications.
Table 13: Top 10 Vulnerabilities in CODELLAMA-13B, and GEMINI-pro 1.0.
@@
1 $\mathcal{DF}$ Dereference failure: NULL pointer 11,741 45.22%
2 $\mathcal{BO}$ Buffer overflow on scanf 5,162 19.88%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 3,528 13.59%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 880 3.39%
5 $\mathcal{DF}$ Dereference failure: forgotten memory 711 2.74%
6 $\mathcal{DF}$ Array bounds violated: upper bound 659 2.54%
7 $\mathcal{AO}$ Arithmetic overflow on add 512 1.97%
8 $\mathcal{DF}$ Array bounds violated: lower bound 465 1.79%
9 $\mathcal{AO}$ Arithmetic overflow on mul 454 1.75%
10 $\mathcal{AO}$ Arithmetic overflow on sub 373 1.44%
GEMINI-pro 1.0
1 $\mathcal{DF}$ Dereference failure: NULL pointer 68,570 57.27%
2 $\mathcal{DF}$ Dereference failure: invalid pointer 13,361 11.16%
3 $\mathcal{BO}$ Buffer overflow on scanf 13,106 10.95%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 4,651 3.88%
5 $\mathcal{DF}$ Dereference failure: forgotten memory 3,433 2.87%
6 $\mathcal{AO}$ Arithmetic overflow on mul 2,450 2.05%
7 $\mathcal{DF}$ Array bounds violated: upper bound 2,154 1.80%
8 $\mathcal{AO}$ Arithmetic overflow on add 1,882 1.57%
9 $\mathcal{DF}$ Array bounds violated: lower bound 1,840 1.54%
10 $\mathcal{AO}$ Arithmetic overflow on sub 1,815 1.52%
Table 14: Top 10 Vulnerabilities in MISTRAL-7B and GPT-4.

### v2

@@ 6 Verification Results
$\bm{\mathcal{VS}}$
$\bm{\mathcal{VU}}$
$\bm{\mathcal{VF}}$
$\bm{\mathcal{ER}}$
Gemma-7B 36 7873678736\,78736 787 $351$ $67.28$ $5.25$ $11.62$ $16.30$ $67.01$ $5.07$
GPT-3.5-turbo 76 1687616876\,16876 168 $616$ $96.79$ $6.07$ $7.29$ $26.09$ $65.07$ $1.55$
Gemini Pro 1.0 38 6953869538\,69538 695 $332$ $98.87$ $4.56$ $9.49$ $24.13$ $63.91$ $2.47$
Falcon2-11B 11 9461194611\,94611 946 $338$ $77.75$ $6.34$ $10.28$ $24.56$ $63.16$ $2.00$
Mistral-7B 9 93499349\,9349 934 $161$ $75.06$ $3.84$ $8.36$ $25.88$ $62.08$ $3.68$
Falcon-180B 68 4636846368\,46368 463 $181$ $71.93$ $4.38$ $6.48$ $28.67$ $62.07$ $2.78$
GPT-4o-mini 38 9213892138\,92138 921 $347$ $103.58$ $3.40$ $4.23$ $36.77$ $57.14$ $1.86$
CodeLlama-13B 11 8381183811\,83811 838 $258$ $83.36$ $4.54$ $15.48$ $29.52$ $52.71$ $2.39$
Llama2-13B 17 7791777917\,77917 779 $207$ $75.51$ $4.12$ $12.36$ $31.78$ $51.30$ $4.56$
FormAI-v2 310 531310531310\,531310 531 $616$ $85.77$ $4.85$ $8.27$ $26.99$ $62.07$ $2.67$
Legend:
Max SLOC: Maximum Source Lines of Code in a sample. Avg SLOC: Average Source Lines of Code per sample. Avg CC: Average Cyclomatic Complexity, which measures the complexity of the code based on the number of linearly independent paths. $\mathcal{VS}$: Verifications Success. $\mathcal{VU}$: Verification Unkown. $\mathcal{VF}$: Verification Failed (vulnerable). $\mathcal{ER}$: Error.
Table 9: Top 10 Violations Across All Categories in FormAI-v2 dataset.
Rank Category Violation Type Count Percentage
1 $\mathcal{DF}$ Dereference failure: NULL pointer 289 548289548289\,548289 548 37.83%
2 $\mathcal{BO}$ Buffer overflow on scanf 214 255214255214\,255214 255 27.99%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 73 8387383873\,83873 838 9.65%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 23 5862358623\,58623 586 3.08%
5 $\mathcal{ABV}$ Array bounds violated: upper bound 23 3802338023\,38023 380 3.05%
6 $\mathcal{DF}$ Dereference failure: forgotten memory 21 1082110821\,10821 108 2.76%
7 $\mathcal{ABV}$ Array bounds violated: lower bound 19 9181991819\,91819 918 2.60%
8 $\mathcal{AO}$ Arithmetic overflow on sub 18 3451834518\,34518 345 2.40%
9 $\mathcal{AO}$ Arithmetic overflow on add 15 9661596615\,96615 966 2.09%
10 $\mathcal{AO}$ Arithmetic overflow on mul 12 4621246212\,46212 462 1.63%
Table 10: Detailed Categorisation of Vulnerabilities in the Entire Dataset
Description Count Percentage
Dereference failures:
 - NULL pointer 289 548289548289\,548289 548 $37.83$%
 - Invalid pointer 73 8387383873\,83873 838 $9.65$%
 - Forgotten memory 21 1082110821\,10821 108 $2.76$%
 - Array bounds violated 23 5862358623\,58623 586 $3.08$%
 - Invalidated dynamic object 3 14531453\,1453 145 $0.41$%
 - Access to object out of bounds 3 22132213\,2213 221 $0.42$%
 - Accessed expired variable pointer 1 22712271\,2271 227 $0.16$%
 - Write access to string constant $913$ $0.12$%
 - Non-dynamic memory $342$ $0.04$%
 - Object accessed with incompatible base type $379$ $0.05$%
 - Oversized field offset $170$ $0.02$%
 - Data object accessed with code type $14$ $0.00$%
Arithmetic overflows:
 - On sub 18 3451834518\,34518 345 $2.40$%
 - On add 15 9661596615\,96615 966 $2.09$%
 - On mul 12 4621246212\,46212 462 $1.63$%
 - IEEE mul 9 67396739\,6739 673 $1.26$%
 - IEEE div 3 52235223\,5223 522 $0.46$%
 - IEEE add 2 37523752\,3752 375 $0.31$%
 - IEEE sub 1 63216321\,6321 632 $0.21$%
 - On div $813$ $0.11$%
 - On shl $972$ $0.13$%
 - On modulus $348$ $0.05$%
 - On neg $155$ $0.02$%
Buffer overflows:
 - On scanf 214 255214255214\,255214 255 $27.99$%
 - On fscanf 8 25282528\,2528 252 $1.08$%
 - On sscanf 4 18441844\,1844 184 $0.55$%
Array bounds violations:
 - Upper bound 23 3802338023\,38023 380 $3.05$%
 - Lower bound 19 9181991819\,91819 918 $2.60$%
 - VLA array size in bytes overflows address space size 4 22242224\,2224 222 $0.55$%
Miscellaneous Vulnerabilities:
 - Division by zero 4 31143114\,3114 311 $0.56$%
 - The pointer to a file object must be a valid argument 1 22512251\,2251 225 $0.16$%
 - Invalid Function argument issues $443$ $0.06$%
 - Same object violation $123$ $0.02$%
 - Operand of free must have zero pointer offset $134$ $0.02$%
During the $500$-second verification time-frame, ESBMC identified 192 757192757192\,757192 757 unique programs with vulnerabilities. In contrast, only 25 6742567425\,67425 674 programs, representing $8.27\%$, were verified as secure. Expanding computational resources may increase the number of programs uncovered from $\mathcal{VU}$, thereby potentially extending the $\mathcal{VF}$ category. These results provide an even better lower bound compared to [9], on what percentage of LLM-generated code is vulnerable. The situation is more concerning than merely stating that 62.07% of the generated files are vulnerable, as a single file can contain multiple vulnerabilities. On average, each file contains $3.97$ vulnerabilities. The total number of property violations detected by ESBMC for the overall dataset is 765 366765366765\,366765 366.
The most common type of vulnerability is related to “Dereference failures” accounting for $54.54\%$ of the cases, predominantly due to NULL pointer issues. This category includes a variety of pointer-related issues, such as invalid pointers, a forgotten memory, and array-bounds violations, among others. “Buffer overflows”, mainly triggered by the scanf function, comprise a significant $27.99$% of the vulnerabilities. This highlights common issues in handling buffer sizes and input functions. “Arithmetic overflows” are also notable, covering various operations like subtraction, addition, multiplication, and division, indicating frequent issues in handling numeric calculations without adequate checks. The table further lists “Array bounds violations” and “Division by zero” as common issues, illustrating challenges in correctly managing arrays and arithmetic operations. A smaller portion of the table covers “Miscellaneous Vulnerabilities” which includes a variety of less frequent but notable issues such as invalid file object pointers and operand violations in memory deallocation. Overall, the data emphasizes the need for robust handling of pointers, buffers, and numeric operations within the source code to mitigate the risk of vulnerabilities.
6.1 General observation about code complexity
NIST defines Cyclomatic Complexity (CC) as “the amount of decision logic in a source code function” and recommends a maximum value of 10 [115]. According to NIST, “higher numbers are bad and lower numbers are good.” As Figure 7 shows, many individual programs generated by Gemma-7B exceed the threshold of 10. While SLOC and CC cannot be used to determine whether code is vulnerable directly, we observed that higher cyclomatic complexity can lead to an increased likelihood of vulnerabilities. Models such as GPT-3.5-turbo, Gemma-7B, and Falcon2-11B, which have high CC, also display the highest rates of verification failures.
As earlier shown in Table 8, the Avg. CC (Average cyclomatic complexity per sample) and Avg. SLOC (Average Source Lines of Code per sample) provide insight into the complexity of the code generated by a certain model. As previously mentioned, if a model produces only non-vulnerable code, it doesn’t necessarily indicate high quality; it could suggest that the generated code is very simple (e.g., generating only “print ’hello world”’ examples). While observing SLOC and CC cannot precisely determine a model’s code quality, it is interesting to observe that GPT-4o-mini, CodeLlama-13B, and Llama2-13B had the least lowest verification failed results and the lowest CC scores.
The analysis of Table 8 shows that GPT-4o-mini does not necessarily generate shorter or simpler code. It produces the longest C programs, with an average SLOC of $103.48$, and has the highest verification unknown score ($36.77$%), indicating that the ESBMC verification process takes longer for GPT-4o-mini samples. In contrast, Gemma-7B generates the shortest average SLOC and also has the lowest verification unknown result ($16.30$%). Additionally, GPT-4o-mini produces code with a lower CC, which implies better maintainability and quality, while Gemma-7B has a much higher average CC.
(a) GPT4o-mini.
(b) Gemma-7B.
Figure 7: Comparison of Cyclomatic Complexity between GPT4o-mini and Gemma-7B.
6.2 Keyword Frequency
Figure 8: 32 C keyword distribution
When assessing vulnerabilities in LLM-generated code, a key question arises: How does LLM-generated code compare to human-written code? If the generated code differs significantly, the dataset may not support meaningful comparisons with real-world code. A practical starting point for this analysis is comparing keyword frequencies. In real-world C/C++ projects, such as those from GitHub and datasets like BigVul, common keywords include ‘if’, ‘return’, ‘struct’, ‘int’, and ‘const’, while less frequent keywords include ‘continue’, ‘default’, and ‘extern’. Significant differences in keyword frequencies between LLM-generated and real-world code would question the dataset’s validity.
To investigate, we used a token-based keyword-counting method to analyze the frequency of 32 C keywords in each LLM-generated subset. Ideally, LLM-generated code should exhibit a similar keyword distribution to real-world code. Figure 8 shows the normalized keyword frequency (occurrences per million lines of code) for various LLM-generated codes, with BigVul as a real-world benchmark. The heatmap reveals that LLM-generated and real-world code have closely matching keyword distributions, likely due to the LLMs being trained on human-written GitHub projects.
While there are slight variations in the distribution between LLMs and BigVul mainly for the less frequent words, LLMs show great similarity on how they handle statements, expressions, and variables in distinct ways. Note, that while all LLM generated codes are fully compilable on our dataset, this is not the case with BigVul samples and other human written code datasets.
@@ 6.3 Vulnerability Ranking
Table 11 (Parts I, II, and III) provides an overview of the top 10 vulnerabilities generated by each model. Note that raw vulnerability counts are not directly comparable due to the differing number of samples produced by each model. To enable a fair comparison across LLMs, the table also includes the percentage representation of each vulnerability. This analysis does not offer a comprehensive review of all identified CWEs but focuses on vulnerabilities explicitly verified by ESBMC.
Buffer overflow vulnerabilities related to scanf are consistently ranked among the top three across all LLM models. The functions scanf, fscanf, and sscanf do not restrict the input size of their respective buffers, creating a risk of buffer overflow. This vulnerability can allow attackers to execute arbitrary code or trigger crashes. As previously mentioned, these issues relate to several CWEs, including CWE-676, CWE-20, and CWE-787. Although buffer overflow is a type of out-of-bounds write, CWE-787 covers a broader range of vulnerabilities. CWE-120 specifically addresses classic buffer overflow scenarios caused by unchecked input sizes during buffer copy operations. While more complex issues like arithmetic overflows and array bounds violations require deeper programming context, simpler issues such as scanf errors should be easier to avoid. However, all tested models consistently exhibit buffer overflow errors with scanf.
Table 11: Top 10 Vulnerabilities in LLM Generated Code - Part I
@@
GPT-3.5-turbo
1 $\mathcal{BO}$ Buffer overflow on scanf 84 2138421384\,21384 213 38.23%
2 $\mathcal{DF}$ Dereference failure: NULL pointer 56 6905669056\,69056 690 25.74%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 20 6172061720\,61720 617 9.36%
4 $\mathcal{DF}$ Dereference failure: forgotten memory 4 63146314\,6314 631 2.10%
5 $\mathcal{DF}$ Array bounds violated: lower bound 8 10281028\,1028 102 3.68%
6 $\mathcal{DF}$ Array bounds violated: upper bound 8 10181018\,1018 101 3.68%
7 $\mathcal{AO}$ Arithmetic overflow on sub 6 62766276\,6276 627 3.01%
8 $\mathcal{DF}$ Dereference failure: array bounds violated 6 53765376\,5376 537 2.97%
9 $\mathcal{AO}$ Arithmetic overflow on add 5 22852285\,2285 228 2.37%
10 $\mathcal{AO}$ Arithmetic overflow on mul 4 28542854\,2854 285 1.95%
@@
1 $\mathcal{BO}$ Buffer overflow on scanf 49 1754917549\,17549 175 34.37%
2 $\mathcal{DF}$ Dereference failure: NULL pointer 42 1774217742\,17742 177 29.48%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 15 7321573215\,73215 732 11.00%
4 $\mathcal{DF}$ Dereference failure: forgotten memory 5 44254425\,4425 442 3.80%
5 $\mathcal{DF}$ Dereference failure: array bounds violated 4 $54$\,5454 545 3.18%
6 $\mathcal{DF}$ Array bounds violated: upper bound 4 31043104\,3104 310 3.01%
7 $\mathcal{AO}$ Arithmetic overflow on sub 3 31533153\,3153 315 2.32%
8 $\mathcal{DF}$ Array bounds violated: lower bound 3 61136113\,6113 611 2.52%
9 $\mathcal{AO}$ Arithmetic overflow on add 2 85828582\,8582 858 2.00%
10 $\mathcal{BO}$ Buffer overflow on fscanf 2 53225322\,5322 532 1.77%
@@
1 $\mathcal{DF}$ Dereference failure: NULL pointer 17 6301763017\,63017 630 54.45%
2 $\mathcal{DF}$ Dereference failure: invalid pointer 3 08930893\,0893 089 9.54%
3 $\mathcal{BO}$ Buffer overflow on scanf 2 77527752\,7752 775 8.57%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 1 61116111\,6111 611 4.98%
5 $\mathcal{DF}$ Dereference failure: forgotten memory 1 25412541\,2541 254 3.87%
6 $\mathcal{AO}$ Arithmetic overflow on add $883$ 2.73%
7 $\mathcal{DF}$ Array bounds violated: upper bound $818$ 2.53%
8 $\mathcal{AO}$ Arithmetic overflow on mul $599$ 1.85%
9 $\mathcal{AO}$ Arithmetic overflow on sub $571$ 1.76%
10 $\mathcal{BO}$ Division by zero 462 1.43%
Gemma-7B
1 $\mathcal{DF}$ Dereference failure: NULL pointer 59 4335943359\,43359 433 60.50%
2 $\mathcal{BO}$ Buffer overflow on scanf 14 9501495014\,95014 950 15.22%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 3 61736173\,6173 617 3.68%
4 $\mathcal{DF}$ Dereference failure: forgotten memory 3 19131913\,1913 191 3.25%
5 $\mathcal{DF}$ Array bounds violated: upper bound 3 37933793\,3793 379 3.44%
6 $\mathcal{DF}$ Array bounds violated: lower bound 2 78427842\,7842 784 2.83%
7 $\mathcal{AO}$ Arithmetic overflow on sub 2 04020402\,0402 040 2.08%
8 $\mathcal{DF}$ Dereference failure: array bounds violated 1 78617861\,7861 786 1.82%
9 $\mathcal{AO}$ Arithmetic overflow on floating-point ieee_mul 1 15211521\,1521 152 1.17%
10 $\mathcal{AO}$ Arithmetic overflow on add 1 30213021\,3021 302 1.33%
Table 11 (Cont.): Top 10 Vulnerabilities in LLM Generated Code - Part II
@@
1 $\mathcal{DF}$ Dereference failure: NULL pointer 11 5461154611\,54611 546 44.75%
2 $\mathcal{BO}$ Buffer overflow on scanf 5 16951695\,1695 169 20.03%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 3 48134813\,4813 481 13.49%
4 $\mathcal{DF}$ Dereference failure: array bounds violated $897$ 3.48%
5 $\mathcal{DF}$ Dereference failure: forgotten memory $695$ 2.69%
6 $\mathcal{DF}$ Array bounds violated: upper bound $683$ 2.65%
7 $\mathcal{AO}$ Arithmetic overflow on add $524$ 2.03%
8 $\mathcal{DF}$ Array bounds violated: lower bound $465$ 1.80%
9 $\mathcal{AO}$ Arithmetic overflow on mul $456$ 1.77%
10 $\mathcal{AO}$ Arithmetic overflow on sub $380$ 1.47%
Gemini Pro 1.0
1 $\mathcal{DF}$ Dereference failure: NULL pointer 65 3766537665\,37665 376 55.95%
2 $\mathcal{DF}$ Dereference failure: invalid pointer 13 2721327213\,27213 272 11.36%
3 $\mathcal{BO}$ Buffer overflow on scanf 12 9481294812\,94812 948 11.08%
4 $\mathcal{DF}$ Dereference failure: array bounds violated 4 25042504\,2504 250 3.64%
5 $\mathcal{DF}$ Dereference failure: forgotten memory 3 34033403\,3403 340 2.86%
6 $\mathcal{AO}$ Arithmetic overflow on mul 2 46624662\,4662 466 2.11%
7 $\mathcal{DF}$ Array bounds violated: upper bound 2 28522852\,2852 285 1.96%
8 $\mathcal{DF}$ Array bounds violated: lower bound 1 95219521\,9521 952 1.67%
9 $\mathcal{AO}$ Arithmetic overflow on sub 1 89918991\,8991 899 1.63%
10 $\mathcal{AO}$ Arithmetic overflow on add 1 89518951\,8951 895 1.62%
Mistral-7B
1 $\mathcal{DF}$ Dereference failure: NULL pointer 6 29462946\,2946 294 33.17%
2 $\mathcal{BO}$ Buffer overflow on scanf 5 12551255\,1255 125 27.01%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 2 46024602\,4602 460 12.97%
4 $\mathcal{DF}$ Dereference failure: array bounds violated $738$ 3.89%
5 $\mathcal{DF}$ Array bounds violated: lower bound $622$ 3.28%
6 $\mathcal{AO}$ Arithmetic overflow on sub $473$ 2.49%
7 $\mathcal{DF}$ Array bounds violated: upper bound $453$ 2.39%
8 $\mathcal{DF}$ Dereference failure: forgotten memory $414$ 2.18%
9 $\mathcal{AO}$ Arithmetic overflow on add $400$ 2.11%
10 $\mathcal{BO}$ Buffer overflow on sscanf $388$ 2.04%
GPT-4o-mini
1 $\mathcal{BO}$ Buffer overflow on scanf 33 3073330733\,30733 307 42.60%
2 $\mathcal{DF}$ Dereference failure: NULL pointer 17 5391753917\,53917 539 22.43%
3 $\mathcal{DF}$ Dereference failure: invalid pointer 7 05570557\,0557 055 9.02%
4 $\mathcal{AO}$ Arithmetic overflow on sub 2 47924792\,4792 479 3.17%
5 $\mathcal{DF}$ Array bounds violated: upper bound 2 27722772\,2772 277 2.91%
6 $\mathcal{AO}$ Arithmetic overflow on add 2 11421142\,1142 114 2.70%
7 $\mathcal{DF}$ Dereference failure: array bounds violated 1 95619561\,9561 956 2.50%
8 $\mathcal{AO}$ Arithmetic overflow on floating-point ieee_mul 1 85718571\,8571 857 2.38%
9 $\mathcal{AO}$ Arithmetic overflow on mul 1 53615361\,5361 536 1.96%
10 $\mathcal{DF}$ Array bounds violated: lower bound 1 39813981\,3981 398 1.79%
Table 11 (Cont.): Top 10 Vulnerabilities in LLM Generated Code - Part III